# EEG Foundation Models: A Critical Review of Current Progress and Future Directions

### v2

@@ 6.1 Benchmarks and Tools
Benchmarking against prevailing feature paradigms: Quantitative EEG research currently employs two ML feature paradigms: expert-crafted EEG features and task-specific supervised features. As such, data-driven self-supervised features, such as those provided by EEG-FMs, remain in their infancy. Therefore, to establish a compelling case for EEG-FM adoption, studies proposing novel EEG-FMs should evaluate against the prevailing feature paradigms. In doing so, statistical tests of superior performance must be presented wherever necessary, and identical subject data splits must be used to prevent the influence of inter-subject EEG variability.
Ranking EEG-FMs in global EEG competitions: Novel EEG-FM methodology can leverage international predictive modeling competitions to establish real-world performance and utility, as those competitions can highlight the most significant empirical challenges or milestones within the field. Such competitions have already been developed for time series [157], electrocardiograms [158], and speech [159] domains. Notably, within the EEG domain, previous competitions have highlighted issues with transfer learning under distribution shifts[160], abnormality detection[161], coma prognosis[162], and seizure detection[163]. The creation of new EEG competitions also incentivizes the public release of expertly curated and annotated EEG datasets that can further boost EEG-FM research.
Frictionless software: Several software modifications and code debugging are required to run the currently available EEG-FMs. The deep technical nature of EEG-FMs and the computing skills required to run these models present a considerable entry barrier to non-technical or non-computational researchers. It will be necessary for future studies to ensure that EEG-FMs can be used as off-the-shelf feature extraction tools, with no technical modifications required. Graphical software for plug-and-play EEG-FM analytics, such as simple binary classifications, can greatly increase real-world EEG-FM adoption, testing, and experimentation. Encouragingly, a mature ecosystem for large-scale EEG analytics already exists in Python[164, 165, 166] and can support the development of frictionless EEG-FM software.
@@ 6.2 Technical Development
Holistic evaluation schemes: Evaluating EEG-FMs based on their task performance on publicly available EEG datasets/tasks is insufficient to measure their real-world readiness. Novel evaluation schemes can be designed to assess the computational costs, semantic EEG-FM embedding quality, transfer efficiency, and robustness to real-world EEG variability and noise sources. The development of such evaluation schemes can leverage the wealth of expert knowledge already available within the EEG domain.
Privacy-preserving learning: EEG datasets required for large-scale EEG-FM pretraining can be siloed due to patient privacy and legal concerns. Research collaborations involving multiple clinical sites can utilize federated learning[167] techniques to train EEG-FMs without requiring centralized data access or sharing. Such techniques may also support the development of personalized patient-specific models.
Multi-modal learning: Latent representations learned during pretraining may benefit from cross-modal supervision from text reports, video recordings, or other biosignals like electrocardiograms. Some of these modalities are jointly recorded with EEGs in many clinical and scientific recording settings. Notably, downstream applications can reap the performance benefits of cross/multi-modal pretraining without requiring those modalities during evaluation.
Reasoning and agentic capabilities: Future EEG-FMs could be fine-tuned to mimic expert reasoning processes, adhere to expert grading criteria[168] or instructions, and autonomously retrieve relevant evidence from knowledge bases. Augmented healthcare workflows could utilize insights from a specialized EEG agent, along with other modality-specific clinical data agents, to inform clinical decisions.
Scaling down task-specific EEG-FMs: FMs may require model parameter scaling to obtain superior pretrained features. However, large EEG-FMs may not be suitable for applications involving streaming EEG data, real-time EEG processing, or deployment in resource-constrained settings, such as brain-computer interfaces, medical devices, and clinical environments. In such contexts, model distillation[169], pruning[170], or quantization[171] techniques can deliver smaller, computationally inexpensive inference-only models that maintain high task performance.
@@ 6.3 Application Development
Cross- and inter-disciplinary collaboration: The technical development of EEG-FMs should involve close collaboration with scientific and clinical domain experts. Such collaborations can productively constrain and inspire novel modeling decisions, identify novel research gaps, establish clear criteria for real-world success or progress, and facilitate EEG-FM experimentation among niche application-specific audiences.
Model cards to guide EEG-FM selection: The availability of multiple EEG-FMs, each purportedly an off-the-shelf tool, poses a practical dilemma for domain experts who must make a principled selection for their research. Comprehensive comparisons of all available options may not be feasible in discovery-based research. The release of model cards[172] that summarize an EEG-FM’s functional strengths and weaknesses, inherent or dominant inductive biases, compute requirements, and evaluation scope and rankings could inform the model selection process.
Generative applications: Existing EEG-FMs have primarily focused on learning latent, low-dimensional encodings of raw EEG data. Future studies could explore generative EEG-FM applications by increasing the representational capacity of the EEG decoder. Generative EEG-FMs may be useful in EEG data wrangling tasks, including data denoising, repairing corrupted segments, and imputing missing data, among others. Controllable EEG generation may be applied towards data augmentation within EEG representation learning pipelines.
@@ 7 Conclusion
The promise of foundation models, in principle, lies in effective and robust feature learning, feature re-usability, and label efficiency. Our critical analysis of the state-of-the-art EEG-FMs indicates that the first-generation EEG-FMs, inspired by their counterparts in the mainstream vision and language domains, have made early, moderate strides in realizing this promise for the EEG domain. However, to develop universal, robust, and general-purpose EEG feature extractors, next-generation EEG-FMs must prioritize substantial scaling efforts, principled and trustworthy self-supervised representation learning, and practically relevant evaluations. In addition to technical modeling, we believe that future research should also pursue the collaborative development of meaningful EEG benchmarks, applications, and model evaluation schemes that can measurably track the real-world readiness and impact of EEG-FMs. With sustained efforts in these directions, EEG-FMs are poised to advance scientific research, brain-computer interfaces, and clinical decision support systems.
@@ 9 Funding
Table 3: Channel-hours were used to represent and compare the data scale at which the EEG-FMs were pretrained. In most cases, channel-hours per dataset were calculated as the product of the number of channels and the recording durations (in hours) obtained from the original studies. Scalp EEG datasets had a fixed channel count across subjects, whereas iEEG datasets, used by BrainBERT and Brant, had a subject-specific count. For Neuro-GPT and EEGFormer, channel-hours were computed from publicly available dataset statistics and were assumed to be applicable to both studies. For BrainWave, the total patch count reported was used to calculate the channel-hours by converting the patch size (1 second) to hours. Channel-hours pooled across all pretraining data sources are highlighted in bold. N/A represents cases where channel-hours could not be determined.
@@
($=$ Number of Channels $\times$ Recording Length)
@@
BrainBERT Brain TreeBank [80] - - 91$\times$(1.83+1.75+0.5) + 100$\times$(2.42+2.42+2.5+3+4.33+1.75+3.42) + 91$\times$(1.92+2.67+93.92) + 135$\times$(2.58+1.33+1.83) + 205$\times$(1.75+1.77) + 152$\times$(1.15+1.68+2.43) + (109$\times$1.5) + (102$\times$2.28) + (72$\times$1.28) + 173$\times$(1.58+2.17) 14,752.68
Neuro-GPT TUEG [83] 20 27,063 541,260 541,260
Brant Private dataset (unknown source) - - (124$\times$235.34)+ (52$\times$82.39)+ (120$\times$393.09)+ (133$\times$137.45)+ (116$\times$214.22)+ (101$\times$386.70) + (67$\times$111.55) + (47$\times$207.42)+ (134$\times$759.80) 281,860.10
BIOT SHHS [85, 76] 2 42,446 84,892
@@
LaBraM BCI Competition IV-1 [102] 59 8.21 484.39
Emobrain [103] 64 4.94 316.16
Grasp/Lift EEG Challenge [104] 32 11.72 375.04
Inria BCI Challenge [105] 56 29.98 1,678.88
EEG Motor Imagery Dataset [106] 64 47.3 3,027.2
Visual Categorization EEG Data [107] 64 34.35 2,198.4
Resting State EEG Data  [108] 64 3.04 194.56
SEED Series [109, 110, 111] 62 166.75 10,338.5
Siena Scalp EEG Database [84] 31 30.47 944.57
SPIS Resting State Dataset [112] 64 0.83 53.12
Brain-Invaders [113] 32 16 512
TUAR [83] 23 92.22 2,121.06
TUEP [83] 21 591.22 12,415.62
TUSZ [83] 21 1,138.53 23,909.13
TUSL [83] 23 20.59 473.57
Multiple data sources [114, 115, 116, 117, 118] 62 342.23 21,218.26 80,260.46
@@
Multiple data sources [114, 115, 116, 117, 118] 62 342.23 21,218.26 545,576.06
@@
BrainWave Multiple datasets - - 3,162,233,694$\times$(1$\div$3,600) 878,398.25

### v3

@@ 6.1 Benchmarks and Tools
Benchmarking against prevailing feature paradigms: Quantitative EEG research currently employs two ML feature paradigms: expert-crafted EEG features and task-specific supervised features. As such, data-driven self-supervised features, such as those provided by EEG-FMs, remain in their infancy. Therefore, to establish a compelling case for EEG-FM adoption, studies proposing novel EEG-FMs should evaluate against the prevailing feature paradigms. In doing so, statistical tests of superior performance must be presented when necessary. Additionally, identical subject-level splits must be used to control for inter-subject variability, and temporal past/future splits must be used to account for the temporal order of events in longer-term EEG data.
Standardized community-specific benchmarks: The primary failure mode of EEG encoders remains their inability to learn robust latent features that generalize across realistic EEG distribution shifts, including those seen across subjects, age groups, acquisition setups, and experimental or clinical conditions, among other factors. However, each EEG user community (e.g., clinicians, brain computer interface users, neuroscientists) may have different practical requirements or definitions of model robustness and generalization. Therefore, future work can design community-specific EEG-FM benchmarks with clear task definitions, realistic performance criteria beyond traditional metrics, real-world datasets, and expert baselines. In doing so, EEG researchers can draw inspiration from existing FM benchmarks in other data domains. Standardized benchmarks tailored to each community’s needs can significantly boost the chances of real-world translation and adoption of EEG-FMs.
Ranking EEG-FMs in global EEG competitions: Novel EEG-FM methodology can leverage international predictive modeling competitions to establish real-world performance and utility, as those competitions can highlight the most significant empirical challenges or milestones within the field. Such competitions have already been developed for time series [wang2024deep], electrocardiograms [wan2025openecg], and speech [arora2024evaluation] domains. Notably, within the EEG domain, previous competitions have highlighted issues with transfer learning under distribution shifts [wei20222021], abnormality detection [hms-harmful-brain-activity-classification], coma prognosis [reyna2023predicting], seizure detection [dan2024szcore], and seizure forecasting [brinkmann2016crowdsourcing].
Frictionless software: Several software modifications and code debugging are required to run the currently available EEG-FMs. The deep technical nature of EEG-FMs and the computing skills required to run these models pose a significant entry barrier for non-technical or non-computational researchers. Future studies will need to ensure that EEG-FMs can be used as off-the-shelf feature-extraction tools, with no technical modifications required. Graphical software for plug-and-play EEG-FM analytics, such as simple binary classifications, can greatly increase real-world EEG-FM adoption, testing, and experimentation. Encouragingly, a mature ecosystem for large-scale EEG analytics already exists in Python [gramfort2013meg, HBM:HBM23730, schiratti2018ensemble] and can support the development of frictionless EEG-FM software.
@@ 6.2 Technical Development
Holistic evaluation schemes: Future studies may benefit from focusing specifically on out-of-distribution evaluations, as they can directly assess the effectiveness of EEG-FM pretraining and performance in realistic, off-the-shelf, and low-label scenarios. Furthermore, evaluating EEG-FMs based on their task performance on a narrow set of commonly available EEG datasets/tasks is insufficient to measure their real-world readiness. Novel evaluation and model ranking schemes can be designed to assess the computational costs, semantic EEG-FM embedding quality, transfer efficiency, and robustness to real-world EEG variability and noise sources.
Privacy-preserving learning: EEG datasets required for large-scale EEG-FM pretraining can be siloed due to patient privacy and legal concerns. Research collaborations involving multiple clinical sites can utilize federated learning [mcmahan2017communication] techniques to train EEG-FMs without requiring centralized data access or sharing. Such techniques may also support the derivation of personalized, i.e., patient-specific, models from EEG-FMs.
Multi-modal learning: Latent representations learned during pretraining may benefit from cross-modal supervision from text reports, video recordings, or other biosignals, such as electrocardiograms. Some of these modalities are jointly recorded with EEGs in many clinical and scientific recording settings. Notably, downstream applications can reap the performance benefits of multi-modal pretraining without requiring those modalities during evaluation.
Reasoning and agentic capabilities: Future EEG-FMs could be fine-tuned to mimic expert reasoning processes, adhere to expert grading criteria [tatum2016american] or instructions, and autonomously retrieve relevant evidence from verified medical literature repositories and knowledge bases. Augmented healthcare workflows could leverage insights from multiple specialized modality-specific agents to inform clinical decisions.
Scaling down task-specific EEG-FMs: FMs may require model parameter scaling to obtain superior pretrained features. However, large EEG-FMs may not be suitable for applications involving streaming EEG data, real-time EEG processing, or deployment in resource-constrained settings, such as brain-computer interfaces, medical devices, and clinical environments. In such contexts, knowledge distillation [hinton2015distilling], pruning [han2015deep], or quantization [gong2014compressing] techniques can deliver smaller, computationally inexpensive models that maintain high task performance.
Timeseries FMs, task semantics, and inductive biases: The relative maturity of general time-series FMs (TS-FMs) can serve as a source of inspiration for EEG-FM development. For example, an understanding of how TS-FMs learn effectively from large, diverse datasets across economic, energy, traffic, climate, and industrial domains may improve EEG-FM scaling. TS-FM performance may be particularly instructive when an EEG task shares underlying semantics with more general time-series tasks, such as forecasting or anomaly detection. There may exist opportunities to ‘specialize’ TS-FMs by injecting EEG-specific biases into data representation, model design, and the self-supervised learning objective. Future EEG-FM studies can assess the utility of such biases in adapting TS-FMs for EEG tasks and the relative merits and weaknesses of TS-FMs.
@@ 6.3 Application Development
Cross- and inter-disciplinary collaboration: The technical development of EEG-FMs should involve close collaboration with scientific and clinical experts. Such collaborations can productively constrain and inspire novel modeling decisions, identify novel research gaps, establish clear criteria for real-world success or progress, and facilitate EEG-FM experimentation among niche application-specific audiences.
Model cards to guide EEG-FM selection: The availability of multiple EEG-FMs, each purportedly an off-the-shelf tool, poses a practical dilemma for domain experts who must make a principled selection for their research. Comprehensive comparisons of all available options may not be feasible in discovery-based research. The release of model cards [mitchell2019model] that summarize an EEG-FM’s functional strengths and weaknesses, inherent or dominant inductive biases, compute requirements, and evaluation scope and rankings could inform the model selection process.
Clinical applications: There is a substantial opportunity for EEG-FMs to augment the clinical EEG review process in both short- and extended-duration monitoring settings. EEG-FMs could speed up review by guiding the reviewer’s attention to noteworthy temporal segments, brain regions, or channels distinct from background activity, especially in longer recordings, such as in epilepsy- or sleep-related applications. Based on these clues, an EEG-FM coupled with a text decoder could summarize the clinically relevant findings as a structured text report. Additionally, EEG-FMs could support case-based clinical reasoning by surfacing historical records with shared EEG phenotypes. Finally, EEG-FMs could improve the performance of spike and seizure detections generated by EEG reviewing software.
Translational hurdles: Several hurdles and considerations prevent the translation and adoption of EEG-FMs in clinical practice. Testing and validation. Prospective pilot testing for specific use cases using site-specific patient data can provide actionable evidence on the feasibility of model deployment and its benefits to clinical stakeholders. The graceful handling of data noise, corruption, and missing values is required for operational success. Interpretability. Interfaces that relate EEG-FM outputs to clinical domain knowledge and patient-specific physiological factors could increase clinician trust and adoption. Reproducibility. Developers must transparently report and release the data sources, data splits, patient characteristics, and computational pipeline configurations that were utilized to build the EEG-FM to maintain reproducibility. Regulatory considerations. EEG-FMs and task-specific fine-tuned models will be treated as safety-critical medical artificial intelligence (AI) systems that must adhere to stringent national or international guidelines. Therefore, clinical users should focus on defining clear intended use, rigorous validation, and controlled deployment. Attention should be given to documenting data provenance, identifying known limitations and expected failure modes, and establishing human-AI teaming protocols that keep clinicians in the loop. Clinical adoption should follow medical-device regulatory pathways, including risk assessment, post-market monitoring, and predefined change-control procedures to manage model updates as data distributions evolve.
@@ 7 Conclusion
The promise of EEG foundation models, in principle, lies in effective and robust feature learning, feature re-usability, and label efficiency. Our critical analysis of ten early EEG-FMs indicates that these efforts, inspired by their counterparts in the mainstream vision and language domains, have made moderate strides in realizing this promise for the EEG domain. However, to develop universal, robust, and general-purpose EEG feature extractors, future EEG-FMs must prioritize substantial scaling efforts, principled and trustworthy self-supervised representation learning, and practically relevant evaluations. In addition to technical modeling, we believe that future research should also pursue the collaborative development of meaningful EEG benchmarks, including standardized datasets, applications, and holistic model evaluation schemes that can measurably track the real-world readiness and impact of EEG-FMs. With sustained efforts in these directions, EEG-FMs are poised to advance scientific research, brain-computer interfaces, and clinical decision support systems.
@@ 9 Funding
10 Data Availability
This review did not generate any new data. A visual summary of the EEG-FMs included in this review and links to their original code repositories will be made available via a publicly accessible repository upon acceptance.
11 Conflicts of Interest
None.
@@ 11 Conflicts of Interest
Sleep Foundation Models
Our literature search yielded two sleep FMs that were excluded from the main review scope. Below, we summarize their salient aspects for interested readers.
SleepFM [thapa2024sleepfm]: Data. SleepFM is a large-scale multi-modal sleep foundation model trained on over 100,000 hours of polysomnography data from 14,000 participants. The modalities include brain activity (EEG), eye activity (EOG), muscle tone (EMG), cardiac signals (EKG), and respiratory signals. The data are minimally preprocessed and segmented into 30-second clips aligned with expert annotation boundaries. Modeling. SleepFM uses three modality-specific 1D EfficientNet-based encoders. The model is pretrained with both pairwise contrastive learning and a novel leave-one-out contrastive strategy that aligns each modality with the joint representation of the others. SleepFM supports flexible inference, allowing downstream classifiers to use any subset of the input modalities. Evaluation. The resulting embeddings are evaluated through cross-modal retrieval, demographic prediction, and sleep-related classifications, including sleep staging and sleep disordered breathing detection. Results show that SleepFM consistently outperforms supervised convolutional baselines. Ablations demonstrate that multi-modal pretraining that includes EEG is superior to single-modality pretraining. Additionally, SleepFM demonstrated strong out-of-distribution generalization to an unseen site with a different channel configuration.
Self-Supervised Transformer Model Training for a Sleep-EEG Foundation Model [ogg2024self]: Data. The model is pretrained on 10,897 sleep sessions from 9,013 individuals sourced from multiple public datasets. A single central EEG channel (C3/C4) is resampled to 100 Hz, normalized, clipped, and segmented into 30-second epochs. Sequences of 101 epochs with 25% overlap are constructed for subsequent modeling. Modeling. Unsupervised k-means clustering (300 clusters) on spectrogram features provides pseudo-labels for self-supervision. The approach follows a HuBERT [hsu2021hubert]-style masked prediction framework, where the model learns to infer hidden k-means label sequences. The architecture consists of a 7-layer 1D convolutional encoder, a projection layer, positional encoding, and a 4-layer transformer encoder (28.5M parameters). Masking is applied in 10 epoch blocks, and the model is trained for 40 epochs. Evaluation. After pretraining, the model is adapted for the tasks of sleep staging, subject identification, and age prediction. The pretrained model achieves strong representation learning and rapid convergence, outperforming supervised baselines in low-label sleep staging. Experiments with frozen embeddings, i.e., linear probing, suggest that the model can generalize across montages and tasks. Additionally, the masking and k-means-based pseudo-labeling process were shown to support generalization to the external Sleep-EDF dataset.
Table 1: Pretraining data volume. Channel-hours were used to represent and compare the data scale at which the EEG foundation models were pretrained. In most cases, channel-hours per dataset were calculated as the product of the number of channels and the recording durations (in hours) obtained from the original studies. Scalp EEG datasets had a fixed channel count across subjects, whereas intracranial EEG datasets used in BrainBERT and Brant had a subject-specific channel count. For Neuro-GPT and EEGFormer, channel-hours were computed from publicly available dataset statistics and were assumed to be applicable to both studies. For BrainBERT, we used the channel-hours reported by the authors. For BrainWave, the total patch count reported was used to calculate the channel-hours by converting the patch size (1 second) to hours. Channel-hours pooled across all pretraining data sources are highlighted in bold. N/A represents cases where channel-hours could not be determined.
@@
(== Number of Channels $\times$ Recording Length)
@@
BrainBERT Brain TreeBank [data:wang2024brain] - - 4,551 4,551
Neuro-GPT TUEG [data:obeid2016temple] 20 27,063 541,260 541,260
[…]
BIOT SHHS [data:quan1997sleep, data:zhang2018national] 2 42,446 84,892
@@
LaBraM BCI Competition IV-1 [data:blankertz2007non] 59 8.21 484.39
Emobrain [data:savran06_einterface] 64 4.94 316.16
Grasp/Lift EEG Challenge [data:luciw2014multi] 32 11.72 375.04
Inria BCI Challenge [data:margaux2012objective] 56 29.98 1,678.88
EEG Motor Imagery Dataset [data:schalk2004bci2000] 64 47.3 3,027.2
Visual Categorization EEG Data [data:trujillo2019mental] 64 34.35 2,198.4
Resting State EEG Data [data:trujillo2017effect] 64 3.04 194.56
SEED Series [data:zheng2015investigating, data:zheng2018emotionmeter, data:liu2022identifying] 62 166.75 10,338.5
Siena Scalp EEG Database [data:detti2020eeg] 31 30.47 944.57
SPIS Resting State Dataset [data:torkamani2020prediction] 64 0.83 53.12
Brain-Invaders [data:korczowski2019brain] 32 16 512
TUAR [data:obeid2016temple] 23 92.22 2,121.06
TUEP [data:obeid2016temple] 21 591.22 12,415.62
TUSZ [data:obeid2016temple] 21 1,138.53 23,909.13
TUSL [data:obeid2016temple] 23 20.59 473.57
Multiple data sources [data:jiang2021discriminating, data:jiang2023multimodal, data:luo2022multimodal, data:li2021discrimination, data:tao2020emotion] 62 342.23 21,218.26 80,260.46
@@
Multiple data sources [data:jiang2021discriminating, data:jiang2023multimodal, data:luo2022multimodal, data:li2021discrimination, data:tao2020emotion] 62 342.23 21,218.26 545,576.06
@@
[…]
Table 2: EEG foundation model evaluation strategies. Various techniques were used to evaluate EEG foundation models: a) zero-shot transfer learning (ZSTL, in blue) with prior task knowledge but no access to data and labels from the target setting; b) few-shot learning (FSL, in green) with no prior task knowledge but access to a few labels from the target setting; c) linear probing (LP, in pink) where a classifier head is adapted using labels from the target setting; and d) fine-tuning (FT, in yellow) where the backbone and head together are adapted using labels from the target setting. Evaluation scores estimated based on the graphics in the paper are denoted using superscript .
In-Distribution Out-of-Distribution
EEG-FM Task LP FT FSL ZSTL LP FT
BrainBERT Brain TreeBank AUC - 0.59 AUC - 0.83
Neuro-GPT BCI Competition IV Dataset 2a Acc. - 44.3 Acc. - 64.5
Brant Private (Pathology Detection) Acc. - 91.17
MAYO (Pathology Detection) Acc. - 89.40
FNUSA (Pathology Detection) Acc. - 83.51
BIOT TUAB AUC - 0.8815 AUC - 0.8739
TUEV W.F1 - 0.7504 W. F1 - 0.7322
CHB-MIT AUC - 0.8761 AUC - 0.8752
IIIC Seizure W. F1 - 0.5737 W. F1 - 0.5878
EEGFormer TUAB AUC - 0.876
TUAR AUC - 0.827 AUC - 0.852
TUSL AUC - 0.657 AUC - 0.679
TUSZ AUC - 0.883
Neonate Dataset AUC - 0.842
LaBraM TUAB AUC - 0.8835 AUC - 0.9162
TUEV B.Acc - 34.61 B.Acc - 66.16
SEED Series Acc. - 41.02
MoBI RMSE - 0.1196
Mentality TUSZ AUC - 0.72
NeuroLM TUAB AUC - 0.7884
TUEV B.Acc - 0.4679
TUSL B.Acc - 0.6845
HMC B.Acc - 0.6188
SEED Series B.Acc - 0.6034
Workload B.Acc - 0.6345
FoME TUEV F1 - 46.00 F1 - 79.00
Sleep-EDFx F1 - 99.10
SEED Series F1 - 57.89
MAYO (Events Classification) F1 - 78.80 F1 - 87.78
MAYO (Pathology Detection) F1 - 95.13
FNUSA (Events Classification) F1 - 77.44 F1 - 85.44
FNUSA (Pathology Detection) F1 - 91.81
BrainWave CHB-MIT AUC - 0.90
Clonic-6 AUC - 0.79
Atonic-5 AUC - 0.77
Absence-16 AUC - 0.90 AUC - 0.96
DRE-Clinical AUC - 0.75 AUC - 0.85
SD-71 AUC - 0.63 AUC - 0.70
ADHD-Adult AUC - 0.90 AUC - 0.96
ADHD-Child AUC - 0.77 AUC - 0.80
Schizophrenia-28 AUC - 0.72 AUC - 0.88
Depression-122 AUC - 0.75 AUC - 0.78
MDD-64 AUC - 0.90 AUC - 0.92
AD-65 AUC - 0.63 AUC - 0.76
MAYO (Events Classification) AUC - 0.81 AUC - 0.93 AUC - 0.97
FNUSA (Events Classification) AUC - 0.78 AUC - 0.91 AUC - 0.91